OpenStudy (anonymous):

3(3y-6)=6y+27 Whats y?

OpenStudy (anonymous):

3y=45 ==> y=15

OpenStudy (anonymous):

it was 3

OpenStudy (anonymous):

9y-18=6y+27 9y=6y+45 3y=45 y=15

OpenStudy (anonymous):

9y-18=6y+27 9y-6y=27+18 3y=45 y=45/3 y=15.........as meeru95 has right answer like me........